OpenStudy (kitkat_6778):

Explain why the quantum number set (2, 1, -2, -½) is not possible for an electron in an atom.

OpenStudy (cuanchi):

(2, 1, -2, -½) n= 1, 2, 3 .... l = from 0 to n-1 ml= from -l to + l ms= =1/2 or -1/2